9/11 Pager Messages
		
			2001-09-11
			03:52:23
			Arch
			[1610567]
			B
			ALPHA
			 From wi at 03:52 AM EDT: PWHC - 1 mail,0 vacation and 0 ADI are bad [CC:snyder gmcmullen psmith]
		 
		
			2001-09-11
			09:14:18
			Arch
			[1610567]
			B
			ALPHA
			 From NOC at 09:14 AM EDT: Ibm84/85
		 
		
			2001-09-11
			09:14:20
			Arch
			[1610567]
			B
			ALPHA
			 is timing out while trying to connect to the remote host for Pepco. [CC:lvl2]
		 
		
			2001-09-11
			09:56:25
			Arch
			[1610567]
			B
			ALPHA
			 NOC TKT at 09:56 AM EDT: TKT 37066 - PSS having problems with Recon file [CC:lvl2]
		 
		
			2001-09-11
			10:21:53
			Arch
			[1610567]
			B
			ALPHA
			 From NOC TKT at 10:22 AM EDT: tk37089 - Tg131 is generating an  - All circuits are busy -  message. [CC:lvl1]
		 
		
			2001-09-11
			10:31:27
			Arch
			[1610567]
			B
			ALPHA
			 From NOC at 10:30 AM EDT: FYI - Current Volume is 48384 calls.  LAN is monitoring WAN.  Sys dev monitoring replication.  NOC is working with Goldcenter and telecom to maximize TGs. [CC:lvl2]
		 
		
			2001-09-11
			10:58:19
			Arch
			[1610567]
			B
			ALPHA
			 From NOC at 10:57 AM EDT: W1 is experiencing WAN connectivity issues with W2 Reno Tulsa. WTC LAN is working the issue. [CC:lvl2]
		 
		
			2001-09-11
			11:02:45
			Arch
			[1610567]
			B
			ALPHA
			 NOC TKT at 11:02 AM EDT: TKT 37091 - Earth Link reporting connectivity problems. [CC:lvl1]
		 
		
			2001-09-11
			11:03:23
			Arch
			[1610567]
			B
			ALPHA
			 From
		 
		
			2001-09-11
			11:03:25
			Arch
			[1610567]
			B
			ALPHA
			 NOC at 11:01 AM EDT: ATT has invoked call gapping all services are currently functioning.  [CC:lvl2]
		 
		
			2001-09-11
			11:12:21
			Arch
			[1610567]
			B
			ALPHA
			 NOC TKT at 11:11 AM EDT: TKT 37085 - Ibm84/85 is timing out while trying to connect to the remote host for Pepco. [CC:l
		 
		
			2001-09-11
			11:12:23
			Arch
			[1610567]
			B
			ALPHA
			 vl2]
		 
		
			2001-09-11
			11:15:00
			Arch
			[1610567]
			B
			ALPHA
			 From NOC at 11:14 AM EDT: Earth Link issue resolved. [CC:lvl1]
		 
		
			2001-09-11
			11:21:17
			Arch
			[1610567]
			B
			ALPHA
			 From NOC at 11:20 AM EDT: Issue is resolved for tg131 [CC:lvl1]
		 
		
			2001-09-11
			11:43:15
			Arch
			[1610567]
			B
			ALPHA
			 NO
		 
		
			2001-09-11
			11:43:17
			Arch
			[1610567]
			B
			ALPHA
			 C TKT at 11:41 AM EDT: TKT 37045 - Baton Rouge BRI for ISDN back up is currently down [CC:lvl2]
		 
		
			2001-09-11
			11:49:30
			Arch
			[1610567]
			B
			ALPHA
			 NOC TKT at 11:49 AM EDT: TKT 37093 - Summa25 is failing transfers for RG47. [CC:
		 
		
			2001-09-11
			11:49:32
			Arch
			[1610567]
			B
			ALPHA
			 lvl1]
		 
		
			2001-09-11
			11:54:26
			Arch
			[1610567]
			B
			ALPHA
			 From NOC at 11:54 AM EDT: Summa25 issue has been resolved. [CC:lvl1]
		 
		
			2001-09-11
			12:14:09
			Arch
			[1610567]
			B
			ALPHA
			 NOC TKT at 12:14 PM EDT: TKT 37046 - Hampton BRI for ISDN back up is currently down [CC:lvl1]
		 
		
			2001-09-11
			12:31:56
			Arch
			[1610567]
			B
			ALPHA
			 From NOC at 12:31 PM EDT: Ibm84/85 Pepco issue has been resolved.  [CC:lvl2]
		 
		
			2001-09-11
			13:12:28
			Arch
			[1610567]
			B
			ALPHA
			 NOC TKT at 01:12 PM EDT: TKT 37096 - Data Transfer
		 
		
			2001-09-11
			13:12:30
			Arch
			[1610567]
			B
			ALPHA
			  on Ibm137 cannot connect to Rockford. [CC:lvl1]
		 
		
			2001-09-11
			13:13:17
			Arch
			[1610567]
			B
			ALPHA
			 From NOC at 01:12 PM EDT: PSS recon issue has been resolved. [CC:lvl2]
		 
		
			2001-09-11
			13:35:24
			Arch
			[1610567]
			B
			ALPHA
			 NOC TKT at 01:35 PM EDT: TKT 37097 - Summa24 is failing transfers for RG47. [CC:lvl1]
		 
		
			2001-09-11
			14:04:26
			Arch
			[1610567]
			B
			ALPHA
			 NOC TKT at 02:04 PM EDT: TKT 37046 - Hampton BRI for ISDN back up is currently down [CC:lvl2]
		 
		
			2001-09-11
			15:36:36
			Arch
			[1610567]
			B
			ALPHA
			 From NOC at 03:37 PM EDT: Summa24 issue has been resolved. [CC:lvl1]
		 
		
			2001-09-11
			16:29:58
			Arch
			[1610567]
			B
			ALPHA
			 From wi at 04:30 PM EDT: During the last 30 mins for PwCoop, EXFC LO transactions were too low.
		 
		
			2001-09-11
			16:56:38
			Arch
			[1610567]
			B
			ALPHA
			 NOC TKT at 04:57 PM EDT: TKT 37104 - CSI/Vesta recharge for Sams/walmart generating timeouts. [CC:lvl
		 
		
			2001-09-11
			16:56:39
			Arch
			[1610567]
			B
			ALPHA
			 1]
		 
		
			2001-09-11
			18:36:50
			Arch
			[1610567]
			B
			ALPHA
			 From NOC at 06:37 PM EDT: CSI/Vesta recharge for Sams/walmart issue has been resolved. [CC:lvl1]
		 
		
			2001-09-11
			19:11:11
			Arch
			[1610567]
			B
			ALPHA
			 NOC TKT at 07:11 PM EDT: TKT 37116 Initial - Transcription reporting frozen records for 900Gen1/ 900Chase and 900 Wach. - Assignee: Murray, Kevin M - Group: team4 [CC:team4]
		 
		
			2001-09-11
			19:14:11
			Arch
			[1610567]
			B
			ALPHA
			 From NOC at 07:14 PM EDT: Please call the NOC for pegasus having issues w/900gen1/900chase and 900 wach. [CC:team4]
		 
		
			2001-09-11
			19:19:59
			Arch
			[1610567]
			B
			ALPHA
			 NOC TKT at 07:20 PM EDT: TKT 37116 - Transcription reporting frozen records for 900Gen1/ 900Chase and 900 Wach. [CC:lvl
		 
		
			2001-09-11
			19:20:01
			Arch
			[1610567]
			B
			ALPHA
			 1]
		 
		
			2001-09-11
			19:27:29
			Arch
			[1610567]
			B
			ALPHA
			 NOC TKT at 07:28 PM EDT: TKT 37116 Transfer. - Transcription reporting frozen records for 900Gen1/ 900Chase and 900 Wach. Assignee: Smith, Petra E - Group: team4 [CC:team4]
		 
		
			2001-09-11
			19:41:09
			Arch
			[1610567]
			B
			ALPHA
			 NOC TKT at 07:41 PM EDT: TKT 37045 - Baton Rouge BRI for ISDN back up is currently down [CC:lvl2]
		 
		
			2001-09-11
			20:04:12
			Arch
			[1610567]
			B
			ALPHA
			 NOC TKT at 08:04 PM EDT: TKT 37046 - Hampton BRI for ISDN back up is currently down [CC:lvl2]
		 
		
			2001-09-11
			20:07:29
			Arch
			[1610567]
			B
			ALPHA
			 NOC TKT at 08:08 PM EDT: TKT 37116 Closed - Transcription reporting frozen records for 900Gen1/ 900Chase and 900 Wach. Assignee: Smith, Petra E - Group: team4 [CC:team4]
		 
		
			2001-09-11
			20:09:31
			Arch
			[1610567]
			B
			ALPHA
			 From NOC at 08:10 PM EDT: Transcription reporting frozen record issue has been resolved. [CC:lvl1]